OpenStudy (anonymous):

Haris and Ali see a 500.0 kg of Twinkies fall off a delivery truck onto an ice covered road. They start pulling on the crate with a combined force of 1010 N and Rita the truck driver is pulling in the opposite direction, exerting 400 N. What is the acceleration of the crate and how long does it take to reach the door to the physics classroom 350m away that Mr.Cooper is holding open.

OpenStudy (anonymous):

f=ma a=f/m m=500 f=1010-400=610N ==>a=610/500 s=vt+1/2at^2 350=1/2*610/500*t^2 you can know t.

OpenStudy (anonymous):

oh ok thx

OpenStudy (anonymous):

m = 500 kg F(1) = 1010 N F(2) = -400 N d = 350 m F(net) = F(1) + F(2) = 1010 - 400 = 610 N \[F = ma\]so a = 610/500 = 1.22 m/s^2 \[d = 1/2at^2\]so t^2 = 350/(0.5*1.22) = 350/0.61 = 573.77 s^2 thus, t = 23.95 s Answer: t = 23.95 s

OpenStudy (anonymous):

Thank you